OpenStudy (anonymous):

12^x-2=20 im doing logarithms

OpenStudy (anonymous):

12^x=20-2 12^x=22 This is a log property \[a^{b}=C, \log_a{C}=b \] I think this will help just plug in

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome

OpenStudy (jdoe0001):

$$ 12^x-2=20\\ 12^x = 22\\ \text{using log cancellation rule of}\\ \color{blue}{log_aa^x =x}\\ log_{12}12^x = log_{12}22 \implies x = log_{12}22\\ \text{using log change of base rule of }\\ \color{blue}{log_ab = \frac{log_c(b)}{log_c(a)}}\\ x = \frac{log_{10} 22}{log_{10}12} $$